OpenStudy (anonymous):

The ratio of two numbers is 7:4. If 8 is added to the larger number and 4 is subtracted from the smaller number, the ratio of the numbers is 5:2. Find the two original numbers. Thank you thank you thank you~ ^^

Directrix (directrix):

ratio of two numbers is 7:4. Let one number be 7x and the other 4x.

Directrix (directrix):

If 8 is added to the larger number and 4 is subtracted from the smaller number, the ratio of the numbers is 5:2. (7x + 8) / (4x - 4) = 5/2 5(4x - 4) = 2(7x + 8) Solve for x and post what you get.

OpenStudy (anonymous):

I got x=6.

Directrix (directrix):

The two original numbers are 7x and the other 4x. So, one of them is 7 times 6 and the other is 4 times 6. So, what are the original numbers?

OpenStudy (anonymous):

42:24.....?

Directrix (directrix):

Yes, those are correct. The two numbers are 42 and 24.

OpenStudy (anonymous):

Yay, thank you ^^

Directrix (directrix):

You're welcome.